Prev Next

SW&IT Meeting Minutes 2020-01-13

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191230
LR20191230 - Bec, Matthieu 72.0
LR20191230 - Holmberg, Carl 56.0
LR20191230 - Colson, Andrew 56.0
LR20191230 - Edmeades, David 0.0
LR20191230 - Sola, Igor 17.0
LR20191230 - Kubanek, Petr 60.0
LR20191230 - Hooper, Stephen 56.0
LR20191230 - Zhang, Yang 40.0
LR20191230 - Zhang, Xianyu 0.0

News from management

Handover steps

  1. sub-system techtalk presentations (Andrew, Petr)
  2. new primary + secondary do self learning, gather information, read wiki
  3. hands-on walk through with primary and secondary (focus on operation, deployment, development if applicable)

Subsystems


Application(s) Primary Secondary
TCS + MCSPU Yang Carl
Argos Carl Matthieu
OAC + AzCam Carl Yang
DIMM Yang Stephen
FACSUM, Weatherstation Matthieu Stephen
ALH/INDI Matthieu Carl
OVMS Stephen Matthieu
particle monitors Stephen Matthieu

Previous Action Items

Owner Action Due Date Completion Date CommentSorted ascending
DE As we're moving more critical systems into VMs running off SSD, test Veam Backup 2020-01-20    
PK Transition plan for overlap 2019-12-30    
AC Schedule techtalk OAC, DIMM, Weather Station 2019-12-30    
DE Summit Veam Backup (soul-telemetry, etc) + bring back VM2 until PEPSI/VM arrives 2020-01-20    
DE Non-compliant VMware account 2020-01-20    
MB flatten out telemetry on the mountain - talk to John 2020-01-13 2020-01-13 copy as-is
IT SHARK-VIS/Forerunner - the temp, the mid and the long term 2020-01-09 2020-01-06 reschedule
SH DX SOUL Telescope Network upgrade to JumboFrames - IT7969 2019-01-16 CleanRM - 2019-01-29 Tel 2020-01-06 SH it's ready already

Open topics

Team reports

PK

  • 2020A build with Yang, deployed & running
  • RELEASING.md, updated DEVELOP.md
  • Malaga AstroRob 2019 article
  • investigated IT7983 and IT7987
  • prepared 2020B build * PR for mirror cover removal * merged Oil level patch
  • vacation next two Mondays

AC

  • Last week
    • DIMM tech talk and follow-up with YZ
    • OAC tech talk prep - comments from PK
    • Updated DIMM, OAC, and ARGOS wiki
  • This week
    • OAC tech talk and follow-up
    • Misc AGW2 alignment tasks with CH and YZ - guider and wfs images, etc.

MB

  • Last week
    • OC, WPC, Management
    • phone interviews
    • DMS streaming, architecture and websocket+mongodb prototyping
    • helping out SOUL/AGW2 media converter ...
    • telemetry shuffling discussion
  • This week
    • DMS, websocket continued, start documentation
    • AI grad student follow up
    • resume and conclude AO data with XZ
    • telemetry shuffling conclusion
    • PMC out-of-order packets follow up

YZ

  • TCS transition study with Petr:
    • TCS 2020A build and release.
    • PCS patch and rotator debug etc.
    • Reviewing TCS notes and documents.
  • DIMM and OAC transition with Andrew. More to talk about on Monday.
  • Some safety training.

SH

  • AGW in the lab, guider in the lab
  • network issues: LBTI, PEPSI and still bad latency
  • Windows 2008 licensing (bookstore).
  • obs3 issues need to look.

next:
  • NFS root works, but breaks after a second. Could be fun, but why is it breaking? Ovirt, openstack, hyperconvergence.
  • Leroy particle monitor on and weather bridge replacement.
  • Work with Leroy on 410s for PEPSI.
  • Replace right monitor in 507 -- gets stuff out of my office.
  • Figure out what I need to move down.

IS

Last week:
  • OT & PIT
    • Microdither bug
    • Resubmission logic in PIT
  • Safety training
  • Setting up new laptop
This week
  • Safety training
  • PIT
    • ongoing work on resubmission logic mostly backend
  • OT
    • Fix LBC standards and MODS bias scripts with dgo and go through the bug list fixing stuff.

CH

  • Last week
    • AHS: TMS-related laser power meter monnit sensor will stream to HDF5, not DD.
    • ARGOS: rebuilt simulator under own account, need to fix error Ice version mismatch.
    • AZCam: unable to launch simulator container
    • OAC: initial wiki pages review.
    • PK TCS draft paper review.
    • Safety class.
    • TCS: IT7957, add patches per MB and PK suggestions.
  • This week
    • AHS: prototype the LBTI INDI driver.
    • ARGOS: rebuild simulator on Jenkins.
    • DMS: get stub temperature page posting requests to its wsgi module.
    • TCS: test patches.

IS

Last week:
  • Time Accounting Tool
    • Adding editing to the Queue Time Accounting tool
    • Adding new backup and poorweather program tabs and change logic to accrue to the general Backup and Poorweather bins in the Downtime tab.

This week:
  • Time Accounting Tool
    • Find a way to reroute in the server and the client the "reporting" part.
    • Export to excel
    • Look into django admin login capabilities for the tool

Topic revision: r10 - 27 Jan 2020, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback